# FSOS-AMC: Few-Shot Open-Set Learning for Automatic Modulation Classification

### v1

@@ III-B Meta-Protype Training
where $x_{(}i,j)^{Q^{+}}$ and $x_{(}i,j)^{Q^{-}}$ are the query samples from the correct and incorrect classes, respectively. $m$ is the margin parameter, and $N_{t}$ is the number of triplets. d is the distance function, given as
@@
and $-d_{L2}$ is the euclidean distance function.
III-C Modular Open-Set Classifier
@@ IV-A Evaluation Metrics
where $N_{tr}$ represents the number of known modulation schemes in the training set, and $N_{te}$ is the number of known modulation schemes in the testing set. Typically, the openness value is between 0 and 1. A higher openness value indicates a more challenging open set recognition problem. $N_{tr}\leq N_{te}$. When $N_{tr}=N_{te}$, the value of $openness$ is 0 and it is closed set recognition.
To evaluate the performance of the proposed FSOS-AMC framework, classification accuracy is used to evaluate the performance of the proposed FSOS-AMC framework for known classes. Moreover, to evaluate the performance of the proposed FSOS-AMC framework for unknown classes, two metrics, AUROC (Area Under the Receiver Operating Characteristic curve), and FRR (False Reject Rate) are used. The larger the AUROC value, the better the performance of the open set recognition, and the smaller the FRR value, the better the performance.
@@ IV-B Experimental Setup
To evaluate the performance of our proposed FSOS-AMC framework, we conduct experiments on the RadioML 2016.10a dataset [9], which is a widely used dataset for AMC. The dataset contains 11 modulation schemes, including 8 digital modulation schemes and 3 analog modulation schemes at varying signal-to-noise ratios from $-20$ dB to $18$ dB with a step of $2$. The digital modulation schemes include 8PSK, BPSK, CPFSK, GFSK, PAM4, QAM16, QAM64, QPSK. The analog modulation schemes include AM-DSB, AM-SSB, and WBFM. The dataset is generated by GNU Radio and is available on the web. By the specifications outlined in [2], we opt for a signal-to-noise ratio (SNR) range of -6 dB to 12 dB with an interval of 2 dB in our experiment, with total $110,000$ samples. The proportion of the training set and testing set is $80\%$ and $20\%$, respectively. To perform open set recognition, we randomly select 6 modulation schemes (AM-DSB, QAM64, CPFSK, GFSK, 8PSK, PAM4) as the training set and all 11 modulation schemes as the testing set, where 5 modulation schemes are unknown during testing. For this setting, the openness value is $15.98\%$. The dataset configuration is shown in Table I, the compared methods and the proposed FSOS-AMC framework are first trained on the 5 known modulation schemes using a 5-way 10-shot learning manner, and then tested on the 11 modulation schemes.
TABLE I: Dataset Configuration for Few-Shot Open-Set AMC Using RadioML 2016.10a Dataset.
@@
Experiments are conducted on a workstation with an AMD Ryzen 9 7900X3D CPU, 64GB RAM, and an NVIDIA GeForce RTX 4070 Ti GPU, and the system is running Ubuntu 22.04. Code is implemented in Python and the PyTorch deep learning framework. To train our proposed FSOS-AMC framework, we employ the Adam optimization algorithm with an initial learning rate of $0.001$ over $50$ training epochs. During each training epoch, $200$ episodes of $5$ modulation schemes with $10$ samples from each modulation scheme are selected.
IV-C Results and Analysis
To demonstrate the performance of the proposed FSOS-AMC framework for known classes, we first evaluate the accuracy of the proposed FSOS-AMC framework and the compared methods on the 5 known modulation schemes. As shown in Fig. 4, the performance comparison of the proposed FSOS-AMC framework with the state-of-the-art methods, including CNN [9], ResNet [10], MSNet [1], and SSwsrNet [2] in terms of accuracy with respect to the SNR. It can be observed that the proposed FSOS-AMC framework outperforms the compared methods in terms of accuracy, achieving the highest accuracy over $95\%$ across all SNR values. Moreover, the proposed FSOS-AMC framework can achieve an accuracy over $90\%$ when the SNR is larger than 00 dB, which is 2 dB more sensitive than the recent SSwsrNet, and MSNet. The recent SSwsrNet and MSNet achieve a similar performance of $93.6\%$, which is about $1.5\%$ lower than the proposed FSOS-AMC framework.
@@ IV-C Results and Analysis
Figure 4: Classification performance comparison of the proposed FSOS-AMC framework with the state-of-the-art methods, including CNN [9], ResNet [10], MSNet [1], and SSwsrNet [2] under various signal-to-noise ratio (SNR).
To demonstrate the performance of the proposed FSOS-AMC framework for unknown classes, we evaluate the AUROC, and FRR of the proposed FSOS-AMC framework and the compared methods on the 5 unknown modulation schemes. As shown in Fig. 5, the performance comparison of the proposed FSOS-AMC framework with the state-of-the-art methods, including CNN [9], ResNet [10], MSNet [1], and SSwsrNet [2] in terms of AUROC, and FRR. It can be seen that the proposed FSOS-AMC framework achieves the highest AUROC of $54.25\%$, which is about $0.8\%$ higher than the recent SSwsrNet. Moreover, the proposed FSOS-AMC framework achieves a lower FRR compared to other methods.
Figure 5: Performance comparison in terms of AUROC and FRR.
To further demonstrate the effectiveness of the proposed FSOS-AMC framework under open-seg scenarios, we present the confusion matrix of the proposed FSOS-AMC framework and the recent SSwsrNet [2] under SNR = 00dB. It can be found that the proposed FSOS-AMC framework can achieve over $95\%$ classification accuracy for known classes and $36\%$ classification accuracy for unknown classes, which is superior to the recent SSwsrNet. Moreover, compared to the performance of the recent SSwsrNet on unknown classes, the proposed FSOS-AMC framework can achieve a gain of $13\%$ classification accuracy under low SNR conditions.
@@
Figure 6: Confusion matrix of the proposed (a) FSOS-AMC framework and (b) SSwsrNet [2] under SNR = 00 dB.
@@ V Conclusion
A novel few-shot open-set automatic modulation classification (FSOS-AMC) framework is proposed in this paper. The FSOS-AMC framework consists of three main parts, a multi-scale attention feature extractor (MSANet), a meta-prototype training strategy, and a modular open-set classifier. The MSANet is exploited to extract the features from the input signal, and the meta-prototype training strategy is utilized to train the feature extractor using a supervised, episode-based methodology. The modular open-set classifier is adopted to classify the testing data into one of the pre-defined known modulations or potential unknown modulations. Simulation results demonstrate that the proposed FSOS-AMC framework can achieve a high classification accuracy for known modulations and unknown modulations. The proposed FSOS-AMC framework outperforms the recent advanced methods in terms of accuracy, AUROC, and FRR under few-shot open-set scenarios.

### v2

@@ III-D Meta-Protype Training
where $x_{i,j}^{Q^{+}}$ and $x_{i,j}^{Q^{-}}$ are the query samples from the correct and incorrect classes, respectively. $m$ is the margin parameter, and $N_{t}$ is the number of triplets. d is the distance function, given as
@@
and $d_{L2}$ is the Euclidean distance function.
III-E Modular Open-Set Classifier
@@ IV-A Evaluation Metrics
where $N_{tr}$ represents the number of known modulation schemes in the training set, and $N_{te}$ is the number of known modulation schemes in the testing set. Typically, the openness value is between 0 and 1. A higher openness value indicates a more challenging open set recognition problem. $N_{tr}\leq N_{te}$. When $N_{tr}=N_{te}$, the value of opennessopenness is 0 and it is closed set recognition.
To evaluate the performance of the proposed FSOS-AMC framework, the classification accuracy is used to evaluate the performance of the proposed FSOS-AMC framework for known classes. Moreoverm, to evaluate the performance of the proposed FSOS-AMC framework for unknown classes, the metric AUROC (Area Under the Receiver Operating Characteristic curve) is used. The larger the AUROC value, the better the performance of the open set recognition.
@@ IV-B Experimental Setup
To evaluate the performance of our proposed FSOS-AMC framework, we conduct experiments on the RadioML 2016.10a dataset [23], which is a widely used dataset for AMC. The dataset contains 11 modulation schemes, including 8 digital modulation schemes and 3 analog modulation schemes at varying signal-to-noise ratios from 20-20dB to 1818dB with a step of $2$. The digital modulation schemes include 8PSK, BPSK, CPFSK, GFSK, PAM4, QAM16, QAM64, QPSK. The analog modulation schemes include AM-DSB, AM-SSB, and WBFM. The dataset is generated by GNU Radio and is available on the website [^1] https://www.deepsig.ai/datasets. By the specifications outlined in [3], we opt for a signal-to-noise ratio (SNR) range of -6dB to 12 dB with an interval of 2 dB in our experiment, with total $110,000$ samples. The proportion of the training set and testing set is $80\%$ and $20\%$, respectively. To perform open set recognition, we randomly select 6 modulation schemes (AM-DSB, QAM64, CPFSK, GFSK, 8PSK, PAM4) as the training set and all 11 modulation schemes as the testing set, where 5 modulation schemes are unknown during testing. For this setting, the openness value is $15.98\%$. The dataset configuration is shown in Table III, the compared methods and the proposed FSOS-AMC framework are first trained on the 5 known modulation schemes using a 5-way 10-shot learning manner, and then tested on the 11 modulation schemes.
To investigate the performance of the proposed FSOS-AMC framework under multipath fading channels, we conduct experiments on the HisarMod2019.1 dataset [34]. The dataset is a diverse signal dataset created to enhance research in wireless communication and modulation classification. It comprises $780,000$ signals across 26 modulation classes from 5 modulation families (analog, FSK, PAM, PSK, and QAM), as shown in Table I. Each signal consists of 1024 I/Q samples, generated using MATLAB 2017a with an oversampling rate of 2 and raised cosine pulse shaping. The dataset covers 20 SNR levels from -20dB to 18dB and simulates 5 different wireless communication channels (ideal, static, Rayleigh, Rician, and Nakagami-m), with equal distribution among these channels. This comprehensive dataset aims to provide a robust resource for researchers, comparable to the RadioML2016.10a dataset, and includes various fading models to represent real-world communication scenarios, as shown in Table II. The dataset configuration is shown in Table IV, the compared methods and the proposed FSOS-AMC framework are first trained on the 13 known modulation schemes using a 5-way 10-shot learning manner, and then tested on the 26 modulation schemes.
TABLE I: 26 Different Modulation Types in Hisarmod2019.1
Modulation Family Modulation Types
Analog AM-DSB, AM-SC, AM-USB, AM-LSB, FM, PM
FSK 2-FSK, 4-FSK, 8-FSK, 16-FSK
PAM 4-PAM, 8-PAM, 16-PAM
PSK BPSK, QPSK, 8-PSK, 16-PSK, 32-PSK, 64-PSK
QAM 4-QAM, 8-QAM, 16-QAM, 32-QAM, 64-QAM, 128-QAM, 256-QAM
TABLE II: Comparison of HisarMod2019.1 [34] and RadioML2016.10a
Feature HisarMod2019.1 RadioML2016.10a
Total Signals 780,000 220,000
Mod. Classes 26 11
Mod. Families Analog, FSK, PAM, PSK, QAM 8 digital, 3 analog
Signals/Mod. 1,500 1,000
Signal Length 1,024 I/Q samples 128 I/Q samples
SNR Levels 20 (-20dB to 18dB) 20 (-20dB to 18dB)
Channels Ideal, Static, Rayleigh, Rician (k=3), Nakagami-m (m=2) Fading channel with AWGN noise
Signals/Channel 300 -
Oversamp. Rate 2 -
Pulse Shaping Raised cosine (0.35) -
Multipath Taps 4 or 6 -
Gen. Tool MATLAB 2017a GNU Radio
TABLE III: Dataset Configuration for Few-Shot Open-Set AMC Using RadioML 2016.10a Dataset.
@@
TABLE IV: Dataset Configuration for Few-Shot Open-Set AMC Using HisarMod2019.1 Dataset.
Training Testing
Classes 64PSK, AM-DSB, FM, 16PSK, BPSK, 8FSK, 2FSK, 4FSK, QPSK, 128QAM, 4PAM, 8PAM, 64QAM 64PSK, AM-DSB, FM, 16PSK, BPSK, 8FSK, 2FSK, 4FSK, QPSK, 128QAM, 4PAM, 8PAM, 64QAM, unknown: {8PSK, 32PSK, 4QAM, 8QAM, 16QAM, 32QAM, 256QAM, 16FSK, 16PAM, AM-DSB-SC, AM-USB, AM-LSB, PM}
Numbers 5-way 10-shot $130,000$ samples
openness $18.35\%$
@@
Experiments are conducted on a workstation with an AMD Ryzen 9 7900X3D CPU, 64GB RAM, and an NVIDIA GeForce RTX 4070 Ti GPU (12GB VRAM), and the system is running Ubuntu 22.04. Code is implemented in Python and the PyTorch deep learning framework [40]. To train our proposed FSOS-AMC framework, we employ the Adam optimization algorithm with an initial learning rate of $0.001$ over $50$ training epochs. During each training epoch, $200$ episodes of $5$ modulation schemes with $10$ samples from each modulation scheme are selected.
IV-C Performance Comparison with State-of-the-Art Methods under Multipath Fading Channels
To demonstrate the effectiveness of the proposed FSOS-AMC framework, we first compare it with the state-of-the-art methods in terms of accuracy, including CNN [23], ResNet [24], MSNet [2], and SSwsrNet [3] using the RadioML 2016.10A dataset [23]. The proposed MS-MSANet model consistently outperforms the other methods across all SNR levels, achieving the highest accuracy of 82.81%. It is followed closely by MSANet-IQ (80.99%) and SSwsrNet (80.55%). MSNet, ResNet, and VGG show lower performance, with VGG having the lowest overall accuracy at 70.97%. All models demonstrate improved accuracy as SNR increases, with the most significant gains observed between -6 dB and 0 dB. The performance curves tend to plateau or show minimal improvements beyond 6 dB SNR for most models. This comparison highlights the robustness and superior performance of the MS-MSANet model in classification tasks under varying noise conditions within the FSOS-AMC framework.
@@ IV-C Performance Comparison with State-of-the-Art Methods under Multipath Fading Channels
Figure 6: Classification performance comparison of the proposed FSOS-AMC framework with the state-of-the-art methods, including VGG [23], ResNet [24], MSNet [2], and SSwsrNet [3] under various signal-to-noise ratios (SNR) using RadioML 2016.10A [23].
To better understand the training process of the proposed FSOS-AMC framework, we compare the training process including the training loss and the test accuracy of the proposed FSOS-AMC framework ((MS-MSANet)) with the state-of-the-art methods, including CNN [23], ResNet [24], MSNet [2], and SSwsrNet [3] in terms of training loss and test accuracy using RadioML 2016.10A [23], as shown in Fig. 7. As can be observed, the proposed FSOS-AMC framework exhibits superior performance with the fastest convergence, lowest loss, and highest accuracy (82.81%). Notably, MSANet-IQ demonstrates strong performance as the second-best model, showing rapid initial improvement in both loss reduction and accuracy gain. It maintains consistent second-place performance throughout training, achieving approximately 80.99% accuracy by the end, and outperforms VGG, ResNet, MSNet, and SSwsrNet across all epochs. With the multi-sequence signal representation, the proposed FSOS-AMC framework can reach the highest test accuracy with about $20$ epochs, while the other models need more epochs to achieve the highest test accuracy.
Figure 7: The comparison of training process of the propsoed FSOS-AMC framework with the state-of-the-art methods, including CNN [23], ResNet [24], MSNet [2], and SSwsrNet [3] in terms of (a) training loss and (b) test accuracy using RadioML 2016.10A [23].
To further validate the effectiveness of the proposed FSOS-AMC framework especially under multipath fading channels, we compare the performance of the proposed FSOS-AMC framework with the state-of-the-art methods, including CNN [23], ResNet [24], MSNet [2], and SSwsrNet [3] under various signal-to-noise ratio (SNR) levels using HisarMod2019.1 [34], as shown in Fig. 8. The proposed FSOS-AMC framework, represented by MS-MSANet and MSANet-IQ, demonstrates superior performance compared to VGG, ResNet, MSNet, and SSwsrNet. MS-MSANet consistently outperforms all other models across all SNR levels, achieving the highest accuracy of 76.72%, followed by MSANet-IQ at 73.84%. All models show improved performance as SNR increases, indicating better classification in less noisy conditions. The experiment, conducted using HisarMod2019.1 under multipath fading channels, provides a realistic scenario for evaluating these models in challenging signal environments. ResNet shows the lowest overall performance among the compared models, while the FSOS-AMC framework demonstrates its effectiveness in handling various noise levels in signal classification tasks.
Figure 8: Classification performance comparison of the proposed FSOS-AMC framework with the state-of-the-art methods, including VGG [23], ResNet [24], MSNet [2], and SSwsrNet [3] under various signal-to-noise ratios (SNRs) using HisarMod2019.1 [34] under multipath fading channels.
Fig. 9 illustrates the training process of the proposed FSOS-AMC framework with the state-of-the-art methods, including VGG [23], ResNet [24], MSNet [2], and SSwsrNet [3] using HisarMod2019.1 [34] under multipath fading channels. In Fig. 9 (a), MS-MSANet demonstrates the lowest and most rapidly decreasing loss curve, indicating superior learning efficiency. Fig. 9 (b) reveals that MS-MSANet achieves and maintains the highest accuracy throughout the training process, reaching approximately 75-80% by the 100th epoch. MSANet-IQ performs second-best in both loss reduction and accuracy gain. The other models (MSNet, VGG, ResNet, and SSwsrNet) show varying degrees of performance but consistently lag behind the top two. This comprehensive comparison highlights the effectiveness of the FSOS-AMC framework, particularly the MS-MSANet variant, in achieving lower training loss and higher test accuracy compared to existing methods in the challenging scenario of multipath fading channels.
Figure 9: The comparison of training process of the proposed FSOS-AMC framework with the state-of-the-art methods, including CNN [23], ResNet [24], MSNet [2], and SSwsrNet [3] in terms of (a) training loss and (b) test accuracy using HisarMod2019.1 [34] under multipath fading channels.
IV-D Performance using Different Signal Representations
To demonstrate the impact of different signal representations on the classification performance of the proposed FSOS-AMC framework, we compare the performance of the proposed FSOS-AMC framework using different signal representations, including IQ signal, amplitude and phase (AP), and power spectral density (PSD) under various signal-to-noise ratios (SNRs) using RadioML 2016.10A [23]. Fig. 10 illustrates the performance of various MSANet configurations across different SNR levels. The MS-MSANet (IQ+AP+PSD) model demonstrates superior performance, achieving the highest overall accuracy of 82.81%. This is followed closely by MSANet-IQ and MSANet-IQ+AP, both attaining accuracies around 81%. The MSANet-AP, MSANet-IQ+PSD, and MSANet-AP+PSD models show comparable performance, with accuracies ranging from 78.19% to 78.85%. Notably, the MSANet-PSD model exhibits significantly lower accuracy at 51.73%. All models display a general trend of improved accuracy as SNR increases, with the most substantial gains observed between -6 dB and 0 dB. The performance curves tend to plateau at higher SNR levels, particularly above 6 dB.
Figure 10: Classification performance comparison of the proposed FSOS-AMC framework using different signal representations, including IQ signal, amplitude and phase, and power spectral density (PSD) under various signal-to-noise ratios (SNRs) using RadioML 2016.10A [23].
Fig. 11 shows the training process of the proposed FSOS-AMC framework using different signal representations, including IQ signal, amplitude and phase, and power spectral density (PSD) in terms of training loss and test accuracy using RadioML 2016.10A [23], where Fig. 11 (a) shows the training loss and Fig. 11 (b) shows the test accuracy. It can be observed that the proposed FSOS-AMC framework using multi-sequence signal representation achieves the lowest training loss and the highest test accuracy, followed by the amplitude and phase representation. The PSD representation shows the highest training loss and the lowest test accuracy.
Figure 11: The comparison of training process of the proposed FSOS-AMC framework using different signal representations, including IQ signal, amplitude and phase, and power spectral density (PSD) in terms of (a) training loss and (b) test accuracy using RadioML 2016.10A [23].
To further validate the effectiveness of the proposed FSOS-AMC framework using different signal representations, we compare the performance of the proposed FSOS-AMC framework using different signal representations, including IQ signal, amplitude and phase, and power spectral density (PSD) under various signal-to-noise ratios (SNRs) using HisarMod2019.1 [34] under multipath fading channels, as shown in Fig. 12. It can be seen that all techniques generally improve in accuracy as SNR increases, with MS-MSANet (with IQ+AP+PSD) and MSANet-IQ+PSD performing best overall. MSANet-PSD notably underperforms compared to the others. The performances tend to converge at higher SNR values, particularly above 8 dB. This visualization is crucial for researchers in signal processing and modulation classification, offering insights into the effectiveness of different signal representation techniques under varying noise conditions.
Figure 12: Classification performance comparison of the proposed FSOS-AMC framework using different signal representations, including IQ signal, amplitude and phase, and power spectral density (PSD) under various signal-to-noise ratio (SNR) using HisarMod2019.1 [34] under multipath fading channels.
IV-E Performance Comparison under Few-Shot Open-Set Scenarios
To evaluate the performance of the proposed FSOS-AMC framework under few-shot open-set scenarios, we compare the classification performance of the proposed FSOS-AMC framework with the state-of-the-art methods, including CNN [23], ResNet [24], MSNet [2], and SSwsrNet [3] under various signal-to-noise ratio (SNR) levels using RadioML 2016.10A [23], as shown in Fig. 13. It can be found that MS-MSANet consistently outperforms other models across most SNR levels, achieving the highest accuracy of 88.58%. MSANet-IQ and SSwsrNet show comparable performance, with accuracies around 87%. MSNet and ResNet demonstrate lower performance, with VGG having the lowest overall accuracy at 78.21%. The proposed FSOS-AMC framework shows superior performance in classifying known and unknown modulation schemes under few-shot open-set scenarios, highlighting its robustness and effectiveness in handling challenging classification tasks.
Figure 13: Classification performance comparison of the proposed FSOS-AMC framework under various signal-to-noise ratios (SNRs) using RadioML 2016.10A [23] under few-shot open-set scenarios.
To further investigate the performance of the proposed FSOS-AMC framework under few-shot open-set scenarios, we compare the classification performance of the proposed FSOS-AMC framework with the state-of-the-art methods, including VGG [23], ResNet [24], MSNet [2], and SSwsrNet [3] under various signal-to-noise ratio (SNR) levels using HisarMod2019.1 [34] under multipath fading channels, as shown in Fig. 14. The proposed FSOS-AMC framework, represented by MS-MSANet, demonstrates superior performance compared to state-of-the-art methods, achieving the highest accuracy of 76.72%. It is followed closely by MSANet-IQ at 73.84%. All models show improved performance as SNR increases, indicating better classification in less noisy conditions. The experiment, conducted using HisarMod2019.1 under multipath fading channels, provides a realistic scenario for evaluating these models in challenging signal environments. ResNet shows the lowest overall performance among the compared models, while the FSOS-AMC framework demonstrates its effectiveness in handling various noise levels in signal classification tasks.
Figure 14: Classification performance comparison of the proposed FSOS-AMC framework under various signal-to-noise ratios (SNRs) using HisarMod2019.1 [34] under multipath fading channels and few-shot open-set scenarios.
For a more comprehensive evaluation of the proposed FSOS-AMC framework under few-shot open-set scenarios, we present the confusion matrix of the proposed FSOS-AMC framework with different signal representations under multipath fading channels using HisarMod2019.1 [34], as shown in Fig. 15. The figure shows the classification performance across unknown modulation classes and known modulation classes, total 14 modulation classes. It can be seen that the proposed FSOS-AMC framework using multi-sequence signal representation achieves a better classification accuracy compared to the single IQ signal representation across all modulation classes except 8PAM. Moreover, for unknown modulation schemes, the proposed FSOS-AMC framework achieves a higher classification accuracy compared to MSANet-IQ with over $70\%$ accuracy (0.7 vs 0.64). The proposed FSOS-AMC framework demonstrates also shows better discrimination between similar modulations, evident in the sharper contrast between diagonal and off-diagonal elements. Notably, both models perform exceptionally well for higher-order modulations such as QAM and PAM, shown in the lower right corner with dark blue squares and minimal off-diagonal confusion.
@@ IV-E Performance Comparison under Few-Shot Open-Set Scenarios
Figure 15: Confusion matrix of the proposed (a) FSOS-AMC framework (MS-MSANet) and (b) MSANet-IQ using HisarMod2019.1 [34] under multipath fading channels.
To evaluate the performance of different models for open set recognition, we adopt AUROC for both RadioML 2016.10a [23] and HisarMod2019.1 [34], as shown in Fig. 16. It can be observed that the proposed FSOS-AMC framework, represented by MS-MSANet, achieves the highest AUROC scores for both datasets, with 95.39% for RadioML 2016.10A and 96.46% for HisarMod2019.1. The MSANet-IQ [1] shows slightly lower performance, with AUROC scores of 94.86% and 95.39% for the two datasets, respectively. The other models, including VGG [23], ResNet [24], MSNet [2], and SSwsrNet [3], exhibit lower AUROC scores, with VGG showing the lowest performance. The proposed FSOS-AMC framework demonstrates superior performance in open set recognition tasks, highlighting its effectiveness in classifying unknown modulation schemes under challenging conditions.
Figure 16: Performance comparison in terms of AUROC using RadioML2016.10a [23] and HisarMod2019.1 [34].
@@ V Conclusion
A novel few-shot open-set automatic modulation classification (FSOS-AMC) framework was proposed in this paper. The FSOS-AMC framework consists of three main parts, a multi-sequence multi-scale attention feature extractor (MS-MSANet), a meta-prototype training strategy, and a modular open-set classifier. The MS-MSANet was exploited to extract the features from the multi-sequence input signal, and the meta-prototype training strategy was utilized to train the feature extractor using a supervised, episode-based methodology. The modular open-set classifier was adopted to classify the testing data into one of the pre-defined known modulations or potential unknown modulations. Simulation results demonstrated that the proposed FSOS-AMC framework can achieve a high classification accuracy for both known modulations and unknown modulations. The proposed FSOS-AMC framework outperforms the recent advanced methods in terms of accuracy, and AUROC under few-shot open-set scenarios.
Despite the promising results of our FSOS-AMC framework, several challenges and opportunities remain for future research. The current framework faces challenges in balancing computational efficiency with classification accuracy for real-time applications, improving performance under extreme channel conditions, and handling dynamic spectrum environments. Future research directions include integrating transformer architectures to enhance feature extraction while reducing computational complexity, exploring meta-learning approaches to handle new modulation schemes with fewer training samples, and developing online learning capabilities for continuous adaptation to changing wireless environments. These challenges and opportunities suggest a rich landscape for future research in automatic modulation classification, particularly as wireless communications continue to evolve with new standards and requirements.